### v1

# FENIX: Enabling In-Network DNN Inference with FPGA-Enhanced Programmable Switches

Abstract
Machine learning (ML) is increasingly used in network data planes for advanced traffic analysis, but existing solutions (such as FlowLens, N3IC, BoS) still struggle to simultaneously achieve low latency, high throughput, and high accuracy. To address these challenges, we present FENIX, a hybrid in-network ML system that performs feature extraction on programmable switch ASICs and deep neural network inference on FPGAs. FENIX introduces a Data Engine that leverages a probabilistic token bucket algorithm to control the sending rate of feature streams, effectively addressing the throughput gap between programmable switch ASICs and FPGAs. In addition, FENIX designs a Model Engine to enable high-accuracy deep neural network inference in the network, overcoming the difficulty of deploying complex models on resource-constrained switch chips. We implement FENIX on a programmable switch platform that integrates a Tofino ASIC and a ZU19EG FPGA directly, and evaluate it on real-world network traffic datasets. Our results show that FENIX achieves microsecond-level inference latency and multi-terabit throughput with low hardware overhead, and delivers over 90% accuracy on mainstream network traffic classification tasks, outperforming SOTA.
1 Introduction
Machine learning (ML) is increasingly transforming networked systems. Recent work demonstrates that both traditional ML techniques—such as decision trees, random forests, and support vector machines—and deep learning (DL) approaches can significantly improve networking tasks, including malicious traffic detection [30, 18, 57, 24] and application traffic classification [10, 57, 51]. These advances have been enabled by the emergence of programmable network devices, including P4 switches [11], NetFPGA [31], and SmartNICs [17]. By integrating data-driven learning models with traditional rule-based mechanisms, these systems enable more adaptive and responsive in-network traffic analysis.
Early intelligent network designs typically split responsibilities between planes: the programmable data plane extracts traffic features, while the control plane performs ML inference. Although this leverages existing programmable infrastructure, it introduces communication latencies—often milliseconds [10, 44]. Such delays are especially problematic for time-sensitive applications like intrusion detection, as analysis may lag behind real-time traffic (see § 7 for details).
To address these latency challenges, researchers have explored moving ML inference closer to the data path. One promising approach is to implement inference directly on programmable devices such as SmartNICs, as demonstrated by systems like N3IC [40]. This offloads computation from endpoint CPUs and significantly reduces inference latency. However, new bottlenecks arise: for example, N3IC achieves a maximum throughput of only 40 Gbps, and even the latest commercial SmartNICs are limited to around 400 Gbps [32]. Such throughput is often insufficient for core networks, where switch ASICs handle 4 to 12 Tbps [22, 21], far exceeding current SmartNIC capabilities.
Recent advances in in-network ML have led researchers to implement ML models directly on programmable switch ASICs. For example, systems such as Leo [24] and NetBeacon [57] deploy decision tree and random forest models on Tofino switches using Match Action Tables (MAT), while BoS [51] demonstrates the feasibility of running binary recurrent neural networks (RNNs) on similar hardware. These approaches can potentially enable line-rate ML inference at significantly higher throughput. However, they are typically subject to stringent computational resource constraints, which limit their effectiveness for complex multi-classification tasks (see Section § 7 for details).
These observations raise several key questions: Why do existing switch ASICs struggle to support more sophisticated ML tasks? Beyond the programmable devices discussed above, what other platforms can effectively balance accuracy and latency for in-network ML? In this paper, we present FENIX (FPGA-Enabled Neural Inference eXecution for network switches), a system designed to achieve low latency, high throughput, and high accuracy for in-network intelligent traffic analysis.
Figure 1: Design sapce in Intelligent Network.
[…]
FENIX consists of two main components: the Data Engine and the Model Engine, which together enable in-network traffic analysis on the data plane. The Data Engine extracts relevant features from high-throughput network traffic, while the Model Engine performs machine learning inference on these features. This modular design enables FENIX to achieve three key goals: low latency by avoiding software processing and enabling direct hardware interfacing, high throughput through efficient feature extraction and communication, and high accuracy by supporting full DNN models with minimal quantization loss.
Contributions. The main contribution of this paper is the design, implementation, and evaluation of FENIX, the first system to leverage FPGA augmentation to extend the capabilities of programmable switches for in-network machine learning. Over six months, we designed and manufactured a custom switch, integrating a programmable Tofino switch chip and a ZU19EG FPGA on Printed Circuit Board (PCB). We evaluate FENIX on challenging tasks, including VPN encrypted traffic classification and malware detection. Experimental results show that FENIX achieves up to 537$\times$ lower inference latency compared to control plane-based approaches, and improves classification accuracy by up to 21% over SOTA. Our prototype shows low hardware resource overhead, indicating that FPGA-enhanced switches are practical for deployment in high-speed network environments.
2 BACKGROUND AND MOTIVATION
This section reviews the evolution of DNN models for networking, examines the limitations of programmable switches, highlights the promise of FPGA-assisted networking, and outlines the integration challenges that motivate our hybrid system design.
DNN Learning Models for Networking. Classification tasks such as malicious traffic detection [30, 18, 57, 24] and application traffic classification [10, 57, 51] are fundamental for network intelligence. Prior work has demonstrated the feasibility of in-network AI inference by deploying lightweight models—such as decision trees [24], random forests [57], XGBoost [57], binarized MLPs [40], and simplified RNNs [51]—on programmable switches and SmartNICs. While these models enable practical deployment, their limited representational power often makes it difficult to capture complex or dynamic network behaviors. In contrast, more complex DNNs offer stronger nonlinear modeling capabilities and can adapt more effectively to evolving network conditions, leading to improved accuracy on diverse networking tasks. As a result, supporting DNN inference in network devices has the potential to further enhance the flexibility and precision of in-network intelligence, especially in scenarios where lightweight models face limitations.
Programmable Network Data Plane Limitations. Despite the promise of in-network DNN inference, the resource constraints of programmable data planes present significant challenges for deploying sophisticated models. PISA-based architectures are effective for basic machine learning tasks [57, 24], but encounter fundamental limitations with complex neural networks. The PISA instruction set supports only simple operations (e.g., addition, subtraction, bit-shifting, and logical operations), and lacks native support for floating-point arithmetic, multiplication, division, and complex conditionals—operations essential for modern DNNs. Hardware resources are also limited: for example, commercial switches such as Barefoot Tofino 1 offer only 12 pipeline stages and constrained memory resources (120 Mbit SRAM, 6.2 Mbit TCAM) [51], which are insufficient for parameter-rich neural networks. Furthermore, restrictions such as atomic register access prevent efficient implementation of iterative computations required by DNNs. Collectively, these constraints make it impractical to support DNN inference on programmable switches while maintaining line-rate performance, underscoring the need for alternative approaches to bring advanced neural network capabilities into the data plane.
FPGA-assisted Networking. FPGAs offer a promising solution to the computational limitations of programmable data planes while preserving high-performance networking capabilities. These reconfigurable platforms can interface directly with programmable switch ASICs, providing microsecond-level latency for DNN-based processing via customized hardware acceleration  [^1] An FPGA can be directly connected to a switching ASIC as an external module, as implemented in our system; this approach also applies to switching ASICs other than Tofino.. FPGAs strike a favorable balance among energy efficiency, computational flexibility, and performance—surpassing CPUs in energy efficiency and providing more versatile analytical capabilities than fixed-function switch ASICs [53]. Their support for partial dynamic reconfiguration enables network operators to update analytical functions without disrupting service [16], offering valuable operational flexibility. In hybrid architectures with programmable switch ASICs, these advantages are particularly valuable: conventional packet processing remains in the high-speed switch pipeline, while computation-intensive neural network inference is offloaded to specialized FPGA hardware. This clear division of labor enables the practical deployment of DNN models even in resource-constrained network environments.
Mismatch between FPGA and Switch. Despite their complementary capabilities, fundamental architectural differences between programmable switches and FPGAs present significant integration challenges. The two components operate in fundamentally different ways: switches process packets in deterministic pipelines with nanosecond precision, whereas FPGAs perform variable-latency neural computations requiring sophisticated synchronization. More critically, there is a substantial throughput gap between these chips: modern programmable switches operate at multi-terabit-per-second rates [21], while even the latest commercial FPGAs sustain only hundreds of Gbps [1]. This order-of-magnitude disparity makes it infeasible to offload all switch traffic directly to FPGAs for processing. As a result, a central challenge in designing FPGA-assisted networking systems is to efficiently extract and route only the most relevant feature information from high-speed network streams to the FPGA for deep processing, while preserving overall throughput and low latency.
Motivation. Current research in-network traffic analysis mainly follows two directions: adapting machine learning models to fit data plane constraints (e.g., transforming decision trees into match-action table representations) [51, 57, 24, 49, 26], or sampling data for offline analysis on servers [10, 55, 25]. However, both approaches inevitably compromise either model accuracy or introduce significant latency, and thus fall short of modern network requirements. The core challenge is to deploy powerful deep neural networks (DNNs) in high-speed network environments. While DNNs excel at traffic classification and anomaly detection, they face a critical bottleneck: programmable switches offer limited computational resources and restricted instruction sets, making complex DNN execution infeasible; meanwhile, hardware accelerators such as FPGAs, though efficient for neural network computation, provide processing bandwidth orders of magnitude lower than the Tbps-level throughput of switches.
Goals. Our research aims to develop FENIX, an intelligent network data plane system that simultaneously achieves three core technical goals:
Low latency: FENIX uses FPGA parallelism and direct interfacing with programmable switching ASICs to support neural network inference with microsecond-level latency, without involving software stacks or system buses. This hardware-based approach minimizes processing overhead and maximizes data throughput. As a result, FENIX ensures real-time responsiveness for latency-sensitive network applications.
High throughput: FENIX achieves high throughput in two aspects. First, programmable switching ASICs process all network traffic at line rate, performing initial feature extraction without affecting forwarding performance. Second, a novel feature control mechanism optimizes the communication channel between switches and FPGAs by dynamically adjusting per-flow sampling rates. This approach maximizes effective bandwidth utilization while respecting hardware constraints.
High accuracy: FENIX implements different DNNs (e.g. CNN, RNN) on FPGAs using fixed-point quantization, rather than simplified or binarized models. This approach maintains accuracy close to offline analysis and enables support for advanced pattern recognition and temporal analysis tasks.
Figure 2: The architecture of FENIX.
3 FENIX Design Overview
Figure 2 shows the architecture of FENIX, which is designed to address two main challenges in integrating deep learning inference with programmable switches. The first challenge is managing the throughput mismatch between programmable switching ASICs and FPGAs under dynamic and bursty network traffic. The second challenge is enabling efficient deployment of deep neural network inference within the constraints of switch-based environments.
To address the throughput mismatch, FENIX uses a Data Engine that manages feature caching and transmission control. The Data Engine consists of a Flow Tracker, Rate Limiter, and Buffer Manager, which together dynamically determine when flows’ features are transmitted to the FPGA. By coordinating these components, the Data Engine adapts to changing traffic conditions and ensures balanced utilization of both the switch ASIC and FPGA.
To support DNN inference on the data path, FENIX employs a Model Engine responsible for both flow identification and model execution. The Model Engine integrates a Vector I/O Processor with a DNN Inference Module, and supports a variety of model structures, including embedding, fully connected, convolutional, and recurrent layers. By synthesizing model parameters and feature dimensions at design time, FENIX efficiently utilizes FPGA resources while maintaining flexibility for diverse neural network architectures.
With this design, FENIX achieves high-throughput, low-latency, and accurate DNN inference for in-network traffic analysis. The following sections detail the design of each module.
Figure 3: Workflow of Flow Tracker.
4 Data Engine
As shown in Figure 2, the Data Engine of FENIX is implemented on programmable switch ASICs and consists of three key modules working in concert. The Flow Tracker maintains complex hash table structures to accurately record state information for each network flow, including flow duration $T_{i}$ and packet count $C_{i}$, which serve as the basis for subsequent analysis. The Rate Limiter uses an enhanced token bucket algorithm to control the transmission frequency of feature information according to the computed probability function $P(T_{i},C_{i})$, effectively addressing throughput mismatch between the switch and FPGA, and dynamically adapting to the current traffic rate. This probabilistic sampling method, based on flow features, helps maintain system stability under high-traffic conditions while ensuring the quality of inputs for model inference. The Buffer Manager employs ring buffer techniques to temporarily store feature data, attaches the data to mirrored packet headers, and efficiently forwards traffic features to the Model Engine for further processing.
4.1 Flow Tracker
Overview. The Flow Tracker is a critical component of the Data Engine, responsible for tracking network flows and making per-flow decisions. As illustrated in Figure 3, it maintains a Flow Info Table in the switch’s Static Random-Access Memory (SRAM), using truncated hash values of five-tuples (source IP, destination IP, source Port, destination Port, and Protocol) as unique flow identifiers. For each flow, the table records several key fields: the flow hash value (hash) for identifying new flow arrivals and handling table collisions; backlog packet count (bklog_n) and backlog timestamp (bklog_t) for tracking intervals between feature transmissions; classification results (class) for storing inference outcomes from the Model Engine; and buffer index (buff_idx), which records the flow’s position in the Buffer Manager’s ring buffer via modulo operation.
Upon packet arrival, the Flow Tracker first determines whether it is an inference packet from Model Engine. If so, it checks whether the packet belongs to a new flow or is the result of a hash collision, and then initializes or updates the corresponding flow entry. For flows with existing classification results, packets are forwarded based on these results. For flows without a classification, a lightweight decision tree implemented on the switch ASIC provides packet-level preliminary inference.
Figure 4: Details in Flow Tracker.
Flow Counting Mechanism. We define the start time of a flow as the instant when its first packet arrives at the switch, and count the number of flows that send packets within each timeout interval $T_{w}$. This statistical approach, which focuses on newly arrived flows within each interval, helps mitigate bias arising from missing the exact start or end times of flows [43]. The mechanism operates as illustrated in Figure 4 (a): the flow counter detects new flows by checking hash registers associated with each flow. Upon detecting a new flow, the counter increments $n$ by 1. At the end of each $T_{w}$ period, both the hash registers and the flow count $n$ are reset and subsequently recalculated by the control plane, ensuring accurate and up-to-date flow statistics.
Buffer Index Update. As depicted in Figure 4 (b), since the data plane cannot perform modulo operations directly, we maintain a separate buffer index (buff_idx) for each flow that increments with each packet and resets to 1 when reaching buffer size. The flow’s packet counter (Pkt_cnt) independently tracks total packets. This design enables a logical ring buffer where the buffer index cycles through available space, ensuring new data replaces old data in a circular manner.
Figure 5: Workflow of Rate Limiter.
4.2 Rate Limiter
Overview. Figure 5 illustrates our novel probabilistic rate limiter design, which integrates a dynamic probability model with the token bucket algorithm to intelligently allocate network traffic resources. This hybrid mechanism is specifically engineered to regulate interactions between the data engine and the model engine, ensuring fair opportunities for flow inference while preventing overload of the model engine.
The rate limiter maintains global traffic statistics within each timing window $T_{w}$, including Flow_cnt (the total number of flows) and Pkt_cnt (the total number of packets processed). Leveraging these statistics, the system achieves several advantages: high-speed flows are more likely to fail when requesting tokens, thereby preserving inference opportunities for lower-rate flows, and by capping the token bucket size to no more than the queue length, the system can effectively absorb traffic bursts without causing excessive queuing or packet drops.
Probability-Based Token Allocation Model. From a theoretical perspective, the rate limiter constructs a probability function $P(T_{i},C_{i})$ based on real-time global traffic statistics—specifically, the global flow count $N$ and the aggregate packet rate $Q$. Here, $T_{i}$ denotes the elapsed time since flow $i$ last transmitted its features, and $C_{i}$ represents the number of packets transmitted by flow $i$ during this period. By considering both $T_{i}$ and $C_{i}$, the system can approximate the instantaneous rate of each flow. For each flow $i$, the token allocation is guided by two key criteria:
Criterion 1: In the idealized scenario where all flows operate at equal rates, each flow should receive tokens at an average interval of $\frac{N}{V}$, where $V$ is the token generation rate.
Criterion 2: In practical situations with heterogeneous flow rates, token allocation should be proportional to each flow’s speed. Thus, the expected token acquisition interval for flow $i$ should approach $\frac{Q}{Q_{i}V}$, where $Q_{i}=\frac{C_{i}}{T_{i}}$ represents the average packet rate of flow $i$.
The token bucket generation rate $V$ is set according to the communication bandwidth $B$ between engines, FPGA frequency $F$, and feature vector width $W$:
$V=\min(F,B/W)$ (1)
Based on these criteria, we propose the following piecewise probability model. The formal proof of its fairness can be found in Appendix A.
$P_{i}(T_{i},C_{i})=\begin{cases}\frac{C_{i}(VT_{i}-N)}{QT_{i}-NC_{i}},&\text{if }T_{i}\in\left[\frac{N}{V},\frac{QT_{i}}{C_{i}V}\right]\,\left(\frac{N}{V}<\frac{QT_{i}}{C_{i}V}\right)\\ \frac{T_{i}(VC_{i}-Q)}{NC_{i}-QT_{i}},&\text{if }T_{i}\in\left[\frac{QT_{i}}{C_{i}V},\frac{N}{V}\right]\,\left(\frac{N}{V}>\frac{QT_{i}}{C_{i}V}\right)\\ 1,&\text{if }QT_{i}=NC_{i}\text{ and }T_{i}\geq\frac{N}{V}\\ 0,&\text{if }QT_{i}=NC_{i}\text{ and }T_{i}<\frac{N}{V}\end{cases}$ (2)
Figure 6: Probability curves of token generation model.
Probability Model Deployment. To intuitively demonstrate our probability model, Figure 6 plots the function curves under representative network settings. The illustrated scenario involves 1000 concurrent flows, with the model engine processing packets at 75 Mpps and the network sustaining a total throughput of 1000 Mpps (approximately 800 Gbps, assuming an average packet size of 100 bytes). Given that the data plane cannot directly compute complex probability expressions, the rate limiter discretizes the model in the control plane by constructing a lookup table: the ranges of $T_{i}$ and $C_{i}$ are uniformly partitioned, enabling efficient mapping of all possible value pairs to probabilities in the range $[0,1]$. Figure 6 presents both the original probability function and its table-based approximation, illustrating that our implementation closely preserves the intended behavior of the model.
Algorithm 1 Token Bucket Algorithm for Rate Limiter
1:if $T_{last}=0$ then
2:     $T_{last}\leftarrow T_{now}$, $gap\leftarrow 0$triangleright Initialize for first packet
3:else
4:     $gap\leftarrow T_{now}-T_{last}$, $T_{last}\leftarrow T_{now}$triangleright Calculate time gap
5:end if
6:$rand\leftarrow$ Random(), $prob\leftarrow$ LookupProbability()
7:$bucket\leftarrow bucket+gap$triangleright Refill tokens
8:if $rand<prob$ then $\triangleright$ Selected for sampling
9:     if $bucket\geq cost$ then
10:         $bucket\leftarrow bucket-cost$triangleright Consume token
11:         SendFeatureVector()
12:     end if
13:end if
Token Bucket Algorithm Implementation. In our data plane design, the rate limiter tracks three essential state variables within a custom timing window $T_{w}$: the number of active flows, the total packet count, and the last packet arrival time $T_{last}$. The details of our probabilistic token bucket mechanism are presented in Algorithm 1. Upon the arrival of each packet, the algorithm first computes the time interval between the current arrival time $T_{now}$ and $T_{last}$ to determine the appropriate token replenishment (lines 1–4). Next, a random number $rand$ is generated and compared against the probability value $prob$ retrieved from the precomputed lookup table (line 5). If the packet is probabilistically selected ($rand<prob$) and the token bucket contains enough tokens, the required token $cost$ is deducted from the bucket and the feature vector is transmitted (lines 6–8). Otherwise, if the selection fails or tokens are insufficient, the algorithm only updates the token count based on the elapsed $gap$ (lines 9–14).
Discussion. The Rate Limiter offers several notable advantages. By probabilistically denying token requests from high-speed flows, it helps reserve transmission opportunities for slower flows, promoting fairness across diverse traffic patterns. Additionally, by capping the token bucket capacity to not exceed the queue length, the system can efficiently accommodate bursty transmissions without risking buffer overflow. Overall, this module enables more equitable and adaptable traffic management, while maintaining stability and robustness within the system.
4.3 Buffer Manager
[…]
Figure 7: Mechanism of Buffer Manager.
Feature Transfer Process. The Buffer Manager operates in close coordination with the Rate Limiter to manage feature transmission. As illustrated in Figure 7, when the Rate Limiter determines that feature information should be exported, the Buffer Manager reads the buff_idx value from the Flow Tracker (e.g., buff_idx=2 as shown in the figure) and extracts the corresponding feature vectors from the flow’s ring buffer in SRAM, assembling them into the packet header. The newest feature, stored in metadata, is appended to the end of this header. During the final Deparser Stage, the Buffer Manager inserts the constructed header into a mirrored packet. These mirrored packets encapsulate both the flow’s five-tuple identification and the sequence of feature vectors, providing the Model Engine with the necessary context for processing.
5 Model Engine
The Model Engine is responsible for performing low-latency inference on traffic features extracted by the Data Engine and is implemented on an FPGA. As illustrated in Figure 8, the Model Engine comprises two main components: a Vector I/O Processor, which manages flow information, and a DNN Inference Module, which executes the model inference. Upon receiving packets from the Data Engine, the Vector I/O Processor separates each packet into a flow identifier and its corresponding feature vector. Flow identifiers are temporarily stored in the Flow Identifier Queue to preserve ordering, while feature vectors are forwarded to the DNN Inference Module for processing.
Once inference is complete, the results are placed in the Output Queue. Each inference result in the Output Queue is paired with the corresponding flow identifier retrieved from the front of the Flow Identifier Queue, forming a new packet with both fields. This packet is then transmitted back to the Data Engine, enabling the programmable switch to execute appropriate follow-up operations on the relevant flow based on the inference result. This approach maintains correct mapping between flows and their inference results throughout processing.
5.1 Vector I/O Processor
The Vector I/O Processor is responsible for parsing incoming network packets to extract both the flow identifier (such as the five-tuple) and the associated feature vector. To maintain the correspondence between flows and their features throughout the inference process, the processor employs a FIFO queue to store flow identifiers. Meanwhile, the DNN Inference Module utilizes asynchronous FIFO queues for both its input and output, enabling seamless data transfer between modules operating under different clock domains. This asynchronous FIFO design not only decouples the timing dependencies between the Vector I/O Processor and the DNN Inference Module, but also improves system robustness and throughput.
During operation, the Vector I/O Processor continuously monitors the status of both the flow identifier FIFO and the output FIFO of the DNN Inference Module. Whenever both queues contain valid entries, the processor simultaneously dequeues the head elements, assembles the flow identifier and inference result into a new packet, and transmits this packet to the Data Engine. The Data Engine then forwards the result to the programmable switch, ensuring timely and accurate flow-level actions based on the inference outcome. For example, when a flow is identified as malicious based on the inference result, the programmable switch can record this status in its flow register. Subsequently, whenever packets belonging to this flow are encountered, the switch can enforce corresponding actions such as rate limiting or traffic isolation, thereby mitigating potential security threats in real time.
Figure 8: Workflow of Rate Limiter.
5.2 DNN Inference Module
The DNN Inference Module performs neural network inference for traffic analysis directly on the FPGA, as illustrated in Figure 8. The core computation uses a systolic array optimized for INT8 operations, enabling efficient matrix-vector multiplications for several common neural network layers. Our implementation supports embedding lookups mapped to LUTs, fully connected (FC) layers, convolution (Conv) layers, and recurrent layers. These layers are composed sequentially according to the model architecture, with all layer parameters and feature dimensions fixed at synthesis time to match application requirements and FPGA resource constraints. Although the implementation is not fully modular, each layer’s computation is mapped onto the same systolic array, and the data path is managed to support the required layer ordering.
Model parameters, including weights and biases for all supported layers, are loaded from the host into on-chip memory via the network interface. During inference, feature vectors are processed in batches: embedding lookups are performed first, followed by the configured sequence of FC, Conv, and recurrent layers, all using offline-quantized INT8 arithmetic. Asynchronous FIFO queues decouple dataflow between layers and enable efficient pipelining. This design provides low-latency inference for network data while making efficient use of FPGA compute and memory resources.
6 Implementation
Experimental Hardware. As shown in Figure 9, our experiments are conducted on a high-performance programmable switch platform that integrates both FPGA and Tofino chips on a single board. The hardware is implemented using a 22-layer high-performance printed circuit board (PCB), utilizing back-drilling technology to balance 100 Gbps high-speed interconnection performance and manufacturing costs. The switching system is equipped with a dedicated power control network, supporting a core operating current of up to 100 A and peak control currents up to 300 A, with precise impedance control to ensure reliable power delivery.
The overall design process included approximately half a month for requirement analysis, one month for schematic design, two months for PCB design, and one and a half months for PCB manufacturing, soldering, and prototype hardware debugging. The interface design and testing between FPGA and Tofino required an additional month, with software driver development and system-level integration carried out in parallel.
Figure 9: Hardware Platform of FENIX: The FPGA-Enhanced Programmable Switch Implementation.
Figure 9 presents a photograph of the hardware prototype, with the primary components highlighted as follows:
  • 1
    Tofino 2 switch chip [22]: Featuring 20 MAU stages, 200 Mbits of SRAM, and 10.3 Mbits of TCAM per pipeline. We configured this chip to efficiently implement our Data Engine modules while balancing resource utilization across pipeline stages. Multiple 100 Gbps port channels were designed to connect the Tofino chip to the FPGA, enabling high-bandwidth data exchange.
  • 2
    Xilinx ZU19EG FPGA chip [1]: Integrates approximately 80 Mbits of on-chip memory and 1,143,450 logic elements. The FPGA resources were partitioned to accommodate the Model Engine’s Vector I/O Processor and Neural Computing Array, while maintaining timing closure at high clock frequencies. The FPGA is directly connected to the Tofino chip through high-speed port channels, enabling low-latency communication.
  • 3
    Control plane CPU: Programmed to manage the Tofino chip, responsible for configuration, monitoring, and real-time probability model adjustments.
  • 4
    Four front-panel FPGA ports: Designed with configurable transceivers to operate in either 10G or 25G modes for external communication. The design process required careful signal integrity analysis and impedance matching.
  • 5
    BMC module: Customized for hardware status monitoring and remote management of the platform, including thermal management for high-performance components.
Experimental Software. Our development process encompassed multiple hardware and software platforms, requiring the integration of a diverse set of technologies and toolchains. For the Tofino chip, we used SDE version 9.8.0 for software development and debugging, while FPGA programming was conducted using Vivado and Vitis HLS 2024.1 [8]. The prototype system comprises: (1) approximately 1,500 lines of P4 code [11] to implement the Data Engine’s three core modules—Flow Tracker, Rate Limiter, and Buffer Manager—on the programmable switch ASICs, which involved precise implementation of hash table maintenance, token bucket algorithms, and feature cache rings; (2) 1,000 lines of HLS code for the Model Engine’s two main components—Vector I/O Processor and DNN Inference Module—on the FPGA, responsible for parallel operations and efficient matrix processing; (3) 1,500 lines of Python code and 1,000 lines of C++ code for model training, testing, and quantization, ensuring efficient model execution under the resource constraints of the hardware; and (4) 300 lines of Python code for calculating probability models in the control plane, enabling intelligent feature transmission decisions and adaptive flow management. Additionally, we developed an 800-line Python simulation program to evaluate and validate FENIX’s performance and scalability under high-throughput and large-scale network scenarios.
Model Training and Quantization. We trained several neural network architectures, including Convolutional Neural Networks (CNNs) and Recurrent Neural Networks (RNNs), for intelligent traffic analysis and classification tasks. Our approach leverages universal and protocol-agnostic traffic features: sequences of raw packet lengths and inter-packet arrival times, which capture the temporal patterns of network flows without requiring any protocol-specific information. For feature extraction, we implemented a software-based sliding window mechanism to process raw PCAP files and generate feature vectors, each consisting of a sequence of packet length and timing data from a flow, paired with the corresponding class label for supervised learning.
Our model architectures generally include normalization layers to standardize input features, followed by either convolutional layers (for CNNs) or recurrent units (for RNNs) to capture temporal dependencies. These are further connected to fully connected layers that map extracted features to output classes, with ReLU activation functions introducing non-linearity. To address class imbalance in-network traffic datasets, we applied both oversampling and undersampling techniques during preprocessing, thereby preventing overfitting to dominant classes and improving the representation of minority classes. This comprehensive approach results in more robust models with higher accuracy and better generalization for identifying diverse traffic types.
After training, we performed quantization using Vitis-AI [7] tools to convert floating-point model parameters to 8-bit integer (INT-8) format for efficient hardware deployment. The quantization process assigns different decimal point positions to different layers, based on their activation distributions, to preserve model accuracy. Our experimental evaluation shows that quantization substantially reduces computation and storage requirements while maintaining classification accuracy, with only negligible performance degradation. The quantized models are thus well-suited for deployment on resource-constrained FPGA platforms, enabling high-throughput, low-latency inference without sacrificing classification quality or robustness.
7 EVALUATION
Our evaluation addresses several key questions: (i) What classification accuracy can FENIX achieve on real traffic datasets, and how does it compare to other methods? (ii) What computational resources are required by FENIX, including both the programmable switching ASICs during the feature extraction phase and the FPGA resources? (iii) How is the performance of FENIX affected under different scales of concurrent flows and throughput? (iv) As an asynchronous hybrid system, does FENIX introduce noticeable latency overhead when using the FPGA for inference?
To answer these questions, we implemented FENIX and other baselines on an FPGA-enhanced programmable switch using P4[11] and HLS [8] (Details in Section § 6), and conducted comparative evaluations. We used publicly available datasets for VPN encrypted traffic and malware detection for testing.
7.1 Methodology
Testbed Setup. We implemented FENIX on an FPGA-enhanced programmable switch (Details in Section § 6) using P4 and HLS, connected to a high-performance server equipped with dual Intel Xeon 5418Y processors and 512GB of memory. The server is equipped with two NVIDIA MCX75310AAS-NEAT 400G network interface cards, with one dedicated to traffic generation and the other to packet reception. The sending NIC generates synthetic traffic and replays pcap files using DPDK (v23.11.3) pktgen [15], while the receiving NIC collects and analyzes the processed packets using the DPDK environment.
Schemes compared. We compared FENIX with the following 9 schemes:
(a) FENIX-CNN-FLOW: We implemented a flow-level CNN with 3 convolutional layers (64, 128, 256 filters) and 2 fully connected layers (512, 256 neurons). Flow accuracy is calculated based on majority voting of packet classifications within each flow.
(b) FENIX-RNN-FLOW: We developed a flow-level RNN with a single custom RNN cell (128 units) and a dense output layer. The model processes packet length and IPD features through embeddings and classifies flows using the final hidden state.
(c) FlowLens [10]: We implemented packet statistics in Tofino using FlowLens’ FMA code with control plane inference. Our XGBoost model used default parameters[48], with collection windows adjusted based on throughput.
(d) FENIX-CNN-PKT: We created a packet-level CNN with the same network architecture as FENIX-CNN-FLOW. This model processes individual packets independently, with accuracy calculated on per-packet classification results.
(e) FENIX-RNN-PKT: We implemented a packet-based RNN with the same network architecture as FENIX-RNN-FLOW. Accuracy is measured on a per-packet basis, evaluating the model’s ability to correctly classify each individual packet.
(f) Netbeacon [57]: We implemented multi-phase tree models. Each phase uses a Random Forest (3 trees, depth 7) matching their configuration.
(g) Leo [24]: We implemented a decision tree (max depth 22, up to 1024 leaf nodes) on switches using packet length extremes and cumulative flow length.
(h) BoS [51]: We implemented the largest variant with a binarized GRU network (9-bit hidden states), 8 GRU units, 6-bit embeddings, and complete feature embedding-GRU-output architecture.
(i) N3IC [40]: We implemented a binary MLP on SmartNIC using flow-level and packet-level features with hidden layers [128, 64, 10]. Due to hardware constraints, we simulated switch-side logic and inference in software.
Tasks. We use the following tasks to evaluate FENIX, as summarized in Table 1: (i) Encrypted Traffic Classification on VPN: This task aims to classify network traffic that has been encrypted by VPNs. We use the ISCXVPN2016 dataset [19]. (ii) Malware Identification: This task distinguishes between traffic generated by benign applications and various types of malware. We use the USTC-TFC2016 dataset [47].
Metrics. We use flow-level macro-F1 (the average of F1-scores across different classes) as the accuracy metric for FlowLens, while for other schemes we report packet-level macro-F1. Additionally, we provide a breakdown of Precision and Recall for each class to enable more detailed performance analysis.
Table 1: Experimental settings.
Dataset (Task) ISCXVPN2016 [19] USTC-TFC [47]
Training Flows 29,295 101,789
Test Flows 7,328 25,455
Number of Classes 7 12
Class Ratio
11:4:13:10:
18:128:1
92:10:4:14:17:23:
105:1:16:132:27:1
Optimizer AdamW AdamW
Learning Rate 0.01 0.005
Table 2: Performance Comparison of Different Methods on Encrypted Traffic Classification and Malware Detection
Class FENIX F-CNN FENIX F-RNN FlowLens [10] FENIX P-CNN FENIX P-RNN NetBeacon [57] Leo [24] BoS [51] N3IC [40]
Encrypted Traffic Classification (ISCXVPN2016 [19])
Chat 0.883/0.852 0.939/0.882 0.862/0.922 0.917/0.836 0.804/0.653 0.627/0.169 0.489/0.353 0.922/0.913 0.533/0.527
Email 0.924/0.834 0.944/0.924 0.888/0.821 0.862/0.882 0.893/0.746 0.230/0.321 0.333/0.019 0.932/0.925 0.349/0.353
File 0.879/0.849 0.923/0.912 0.860/0.889 0.886/0.797 0.889/0.844 0.861/0.701 0.815/0.720 0.915/0.922 0.820/0.848
P2P 0.977/0.963 0.976/0.988 0.923/0.913 0.911/0.914 0.932/0.947 0.861/0.908 0.853/0.867 0.925/0.917 0.905/0.892
Stream 0.902/0.968 0.919/0.973 0.966/0.959 0.877/0.965 0.894/0.969 0.890/0.976 0.850/0.944 0.916/0.908 0.886/0.938
Voip 0.989/0.995 0.992/0.996 0.998/0.995 0.999/0.998 0.999/0.999 0.986/0.993 0.994/0.997 0.829/0.723 0.994/0.993
Web 0.803/0.662 0.793/0.625 0.700/0.525 0.800/0.861 0.869/0.821 0.860/0.405 0.784/0.046 0.729/0.623 0.856/0.524
Macro-F1 0.890 0.912 0.870 0.892 0.873 0.658 0.578 0.863 0.738
Malware Detection (USTC-TFC [47])
Cridex 0.999/1.000 0.999/1.000 0.999/0.998 0.999/1.000 0.996/1.000 0.933/0.997 0.984/0.995 0.983/0.996 0.866/0.861
FTP 0.999/0.998 0.999/0.998 0.999/0.998 0.997/1.000 0.993/1.000 0.993/0.485 0.928/0.986 0.986/0.791 0.848/0.853
Geodo 0.979/0.881 0.984/0.891 0.945/0.905 0.932/0.343 0.786/0.424 0.899/0.524 0.369/0.188 0.934/0.615 0.857/0.851
Htbot 0.962/0.987 0.959/0.989 0.964/0.984 0.932/0.986 0.938/0.982 0.830/0.782 0.897/0.935 0.919/0.937 0.871/0.867
Neris 0.921/0.594 0.888/0.732 0.902/0.817 0.766/0.473 0.736/0.575 0.665/0.469 0.584/0.453 0.761/0.610 0.852/0.847
Nsis-ay 0.985/0.970 0.971/0.982 0.985/0.986 0.959/0.968 0.962/0.972 0.984/0.912 0.918/0.960 0.963/0.968 0.860/0.855
Warcraft 0.998/0.999 0.996/0.994 0.995/0.993 1.000/1.000 0.999/1.000 0.890/1.000 0.995/0.997 0.956/0.998 0.855/0.859
Zeus 0.932/0.945 0.955/0.863 0.971/0.932 0.962/0.978 0.978/0.958 0.823/0.260 0.895/0.791 0.976/0.951 0.865/0.860
Virut 0.671/0.941 0.760/0.893 0.827/0.908 0.723/0.863 0.763/0.836 0.571/0.815 0.691/0.692 0.744/0.844 0.859/0.855
Weibo 0.684/0.822 0.702/0.927 0.745/0.804 0.655/0.821 0.685/0.817 0.649/0.789 0.653/0.708 0.652/0.837 0.862/0.859
Shifu 0.999/0.966 0.997/0.995 0.982/0.914 0.947/0.829 0.965/0.767 0.194/0.302 0.610/0.593 0.835/0.501 0.862/0.859
SMB 0.700/0.522 0.845/0.504 0.726/0.654 0.632/0.416 0.665/0.492 0.607/0.434 0.555/0.491 0.651/0.406 0.862/0.859
Macro-F1 0.887 0.901 0.914 0.907 0.838 0.670 0.741 0.814 0.858
Note: The four FENIX columns represent models using flow-level statistics with CNN (FENIX F-CNN) and RNN (FENIX F-RNN), as well as packet-level statistics with CNN (FENIX P-CNN) and RNN (FENIX P-RNN).
7.2 Classification Accuracy on Real Datasets
Table 2 presents a comprehensive comparison of accuracy and recall performance of different methods on encrypted traffic classification and malware detection tasks, evaluated at both the flow and packet levels. Across all categories in both tasks, FENIX consistently demonstrates superior accuracy compared to other baselines. At the flow level, the macro-F1 score of FENIX reaches 0.890, which is very close to that of FlowLens (0.870), indicating that FENIX can achieve state-of-the-art detection performance while retaining the advantages of a neural network-based approach.
When evaluated at the packet level, FENIX outperforms all other compared methods across both tasks. On the malware detection benchmark, FENIX achieves a macro-F1 score of 0.907, greatly surpassing tree-based approaches such as NetBeacon (0.670) and showing a clear margin over advanced neural baselines like N3IC (0.858). Compared to other neural models, including Leo (0.741) and BoS (0.814), FENIX also demonstrates consistently higher accuracy.
This performance gap becomes especially pronounced in challenging multiclass scenarios, where traditional methods often suffer from error propagation or limited packet-level representational capacity. While BoS and Leo achieve moderate results, they still lag behind FENIX, highlighting the benefits of high-precision inference enabled by the FENIX architecture. Notably, BoS is constrained by model binarization and the limited size of each component that switching ASICs can accommodate, which reduces its accuracy. Tree-based methods like NetBeacon can only update predictions at discrete points, further limiting performance on fine-grained, packet-level tasks. Although FENIX does not perform continuous inference for every packet, its higher per-inference accuracy ensures better overall packet-level accuracy. This advantage is particularly evident in complex multiclass settings, where the CNN-based FENIX model achieves the highest classification performance among all tested methods.
Table 3: P4 Systems Resource Overhead Comparison
System SRAM TCAM Bus Stage
FENIX 12.9% 4.4% 3.5% 9
FlowLens [10] 34.2% 0.0% 2.4% 9
BoS [51] 26.3% 6.3% 8.6% 12
Leo [24] 26.9% 9.0% 5.2% 12
NetBeacon [57] 11.6% 18.8% 6.4% 12
7.3 Hardware Resource Utilization
Table 3 compares the hardware resource overhead of several representative P4 systems on programmable switches. The results show that FENIX achieves relatively low resource usage across multiple dimensions. For instance, its SRAM consumption (12.9%) is much lower than FlowLens (34.2%), indicating more efficient memory management. In terms of TCAM usage, FENIX also stands out, with only 4.4% overhead compared to NetBeacon’s 18.8%, highlighting that FENIX avoids the trade-off of reducing one resource at the cost of sharply increasing another. Beyond memory, FENIX maintains efficient use of bus bandwidth and pipeline stages, remaining below or comparable to most other systems. This balanced and predictable resource profile is enabled by the decoupled DataEngine design, which keeps FENIX ’s overhead stable across different tasks and workloads.
Figure 10: Scaling test of FENIX
Table 4 summarizes the FPGA resource utilization, where percentages represent the proportion of total available resources consumed by each module. The results show that LUTs and FFs are the primary resources consumed, particularly in the core computational components of the CNN and RNN modules. For example, the overall CNN module uses 38.4% of LUTs and 33.8% of FFs. While BRAM and DSP usage is higher in the overall modules (up to 7.1% and 8.1% for CNN, and 6.3% and 4.6% for RNN, respectively), it remains minimal within individual submodules, where most values are below 4%. This indicates that our design primarily leverages combinational and sequential logic to efficiently implement neural network inference on the FPGA, while memory and DSP consumption remains moderate. Overall, the resource utilization is low, leaving ample headroom for further optimization and larger deployments.
Table 4: Neural Network Resource Utilization (%)
Module LUT FF BRAM DSP
CNN (overall) 38.4% 33.8% 7.1% 8.1%
Embedding 4.2% 5.1% 0.5% 0.0%
Convolutional 25.6% 19.7% 4.0% 5.7%
FC 8.6% 9.0% 2.6% 2.4%
RNN (overall) 25.6% 31.2% 6.3% 4.6%
Embedding 4.2% 5.1% 0.5% 0.0%
Recurrent 15.8% 18.7% 3.6% 2.4%
FC 8.6% 9.2% 2.2% 2.2%
Vector I/O 6.0% 4.8% 0.3% 0.0%
7.4 Flow Count and Throughput Scalability
We conduct stress tests on FENIX under high-concurrency and high-throughput scenarios. Since all the original network trace files were collected in low-bandwidth networks (e.g., tens of Mbps), we generate high-throughput trace files by concurrently packaging a large number of flows (ensuring each flow has a unique identifier) and reassigning new timestamps to these packets  [^2] We write the original timestamp into the packet header to handle changes caused by flow acceleration.. We then replay these pcap files on our testbed to generate actual network traffic.
Figure 10 presents the results of the scalability test, where we progressively increase the flow concurrency until the physical bandwidth of the traffic generator’s NIC is fully saturated. The results demonstrate that FENIX can easily handle this level of scale, as its macro-F1 scores remain nearly identical to those in Table 2.
To evaluate FENIX at even larger scales, we build a simulator to emulate the entire workflow of FENIX. By replicating the experimental settings of Table 2 and Figure 10, we validate the accuracy of the simulator, finding that the simulation results are almost the same as those from our testbed. Subsequently, we use the simulator to explore significantly larger scales, progressively increasing the flow concurrency to tens of thousands of flows per second and the aggregate throughput to the Tbps level. As shown in Figure 10, the macro-F1 score of FENIX experiences only a minor decrease, with a reduction of about 13.2% at the largest scale.
Figure 11: Average Latency Compare of FENIX and FlowLens.
7.5 Latency Microbenchmark.
We conduct a comprehensive latency breakdown comparing the control-plane inference approach (FlowLens [10]) with our proposed FENIX, as summarized in Figure 11. To capture microsecond-scale delays, we use an RTT-based method: internal transmission is measured via PCB interconnects, and external transmission via optical modules. Inference latency reflects the total model execution time for each system.
[…]
Most importantly, FENIX reduces inference latency by nearly three orders of magnitude compared to FlowLens. While FlowLens requires over $1,000\mu$s for each inference, FENIX completes inference in just $1.2\mu$s on average. This improvement comes from in-network acceleration on FPGA, which offloads model execution from the host CPU and enables efficient, microsecond-scale inference.
8 Discussion
Hardware Adaptability Beyond Tofino. FENIX is prototyped using a Tofino 2 switch and Xilinx ZU19EG FPGA, but its architecture is fundamentally portable. The discontinuation of Intel’s Tofino line [23] does not threaten the broader applicability of FENIX, as the programmable switch ecosystem remains rich, including Juniper Trio [52], Broadcom Trident [20], and RMT-based research designs such as dRMT [13] and Menshen [46]. Industry trends—exemplified by Broadcom Trident 5’s NetGNT, which integrates neural traffic analysis into switching silicon—indicate that intelligent data planes are becoming mainstream, though open programming interfaces for such capabilities remain limited. The hardware and software co-design of FENIX anticipates industry trends and provides a platform that can be easily adapted to new generations of programmable switches.
Prototyping for ASIC Evolution. While custom ASICs like Taurus [44] have set a precedent for per-packet ML inference in data planes—often validated using novel chips like Plasticine [33]—barriers to ASIC adoption remain high due to inflexible design and long development cycles. FENIX addresses this gap by providing a commercial, reprogrammable prototype that achieves high throughput and low latency at switch nodes, serving as both a practical solution and a reference for future ASIC designers. By demonstrating how DNN inference can be efficiently distributed across programmable switches and FPGAs, FENIX charts a pragmatic path toward intelligent data planes in production settings.
Automation and Future Directions. Recent frameworks such as Homunculus [45] highlight the growing need for automated ML pipelines and model deployment in data plane environments. While FENIX currently requires expert-driven integration when migrating to new tasks or hardware, its modular Model Engine design is well-suited for future automation. Extending FENIX with high-level, task-agnostic abstractions and automated toolchains for FPGA and switch targets would further lower deployment barriers and accelerate its adoption across diverse operational scenarios.
9 Related Work
ML for Network Traffic Analysis. Machine learning has become a core tool for network traffic analysis, powering solutions for malicious traffic detection [30, 6, 18, 34], website fingerprinting [36, 14, 38], and fine-grained traffic classification [4, 39]. The field is rapidly advancing toward more sophisticated models and tasks, including traffic prediction and encrypted flow analysis [28, 54, 56]. However, deploying these models at line rate in the data plane remains challenging due to stringent hardware constraints—motivating innovations in model compression, quantization, and hardware-aware co-design.
Data Plane Device Roles and Limitations. Modern network infrastructures rely on programmable switches, SmartNICs, and FPGAs—each with distinct strengths. SmartNICs efficiently offload compute from servers for functions such as rate limiting, traffic analysis, and protocol acceleration [35, 40, 41, 9]. Programmable switches provide unparalleled throughput and enable in-network tasks such as DDoS defense and real-time classification [5, 29, 50, 27, 51, 57, 26, 2, 24, 49, 12, 3]. FPGAs serve as both prototyping platforms and as flexible data plane accelerators, often complementing the rigid resource budgets of ASIC-based switches [42, 44, 16, 37, 53].
Positioning of FENIX. Prior work has either focused on deploying quantized or highly compressed ML models in programmable switches—often sacrificing model complexity and accuracy—or has relied on executing high-precision models at the network edge, typically with limited throughput and incomplete visibility. By contrast, FENIX explores a new point in the design space: it leverages the tight integration of FPGA and programmable switch hardware at central network nodes to support efficient, high-throughput DNN inference in the data plane, achieving a practical trade-off between inference accuracy, adaptability, and deployment feasibility. This hybrid approach demonstrates that advanced neural inference capabilities can be meaningfully integrated into switch-centric architectures without requiring full line-rate DNN execution, offering a pragmatic path toward more intelligent and flexible network data planes, and informing future ASIC and automation developments.
10 CONCLUSION
In this paper, we present FENIX, a hybrid in-network machine learning system designed to enable fast and accurate traffic analysis directly in the data plane. FENIX splits the workload by performing feature extraction on programmable switches and delegating DNN inference to FPGAs, allowing the system to achieve both high throughput and low latency without sacrificing accuracy. To address the bandwidth gap between switches and FPGAs, FENIX introduces a token bucket-based mechanism to efficiently regulate feature transmission. Our implementation and evaluation on real-world traffic show that FENIX delivers microsecond-level inference latency, multi-terabit throughput, and over 90% classification accuracy with minimal hardware overhead. Our results serve as a compelling proof-of-concept for the viability of future intelligent data-plane switches.
Appendix A Problem Formulation
Variable Description
$V$ Token generation rate
$Q$ Global packet rate
$N$ Number of flows
$T_{i}$ Duration of backlogged packets
$C_{i}$ Number of backlogged packets
$Q_{i}$ Current flow packet rate
$P_{i}$ Probability of generating feature vector
Table 5: Variables of token generation probability model.
As described in the main text, our token allocation model is based on two fundamental criteria:
Criterion 1: If all flows have the same rate, to achieve fair allocation, each flow’s transmission period should be $\frac{N}{V}$.
Criterion 2: In practical scenarios where flow rates differ, tokens should be allocated proportionally to flow rates, resulting in a transmission period of $\frac{Q}{Q_{i}V}$ for each flow, where $Q_{i}=\frac{C_{i}}{T_{i}}$ represents the current flow’s packet rate.
Based on these criteria, we develop a probability model that determines when a flow should send feature information to the FPGA. The model ensures that flows of different rates have appropriate opportunities to transmit, while maintaining overall system stability.
For a flow with packet rate $Q_{i}=\frac{C_{i}}{T_{i}}$, we establish a time-dependent linear probability model that varies according to the relationship between $\frac{N}{V}$ and $\frac{Q}{Q_{i}V}$. The mathematical expression for this probability model is as follows:
$P_{i}(T_{i},C_{i})=\begin{cases}\frac{C_{i}(VT_{i}-N)}{QT_{i}-NC_{i}},&\text{if }T_{i}\in\left[\frac{N}{V},\frac{Q}{Q_{i}V}\right],\left(\frac{N}{V}<\frac{Q}{Q_{i}V}\right)\\ \frac{T_{i}(VC_{i}-Q)}{NC_{i}-QT_{i}},&\text{if }T_{i}\in\left[\frac{Q}{Q_{i}V},\frac{N}{V}\right],\left(\frac{N}{V}>\frac{Q}{Q_{i}V}\right)\\ 1,&\text{if }QT_{i}=NC_{i}\text{ and }T_{i}\geq\frac{N}{V}\\ 0,&\text{if }QT_{i}=NC_{i}\text{ and }T_{i}<\frac{N}{V}\end{cases}$ (3)
To verify this model theoretically, we analyze the expected transmission periods. For any flow $i$, the probability of sending feature information over time can be represented as a cumulative distribution function:
$P(t)=\frac{t-t_{s}}{t_{e}-t_{s}}\quad(t\in[t_{s},t_{e}])$ (4)
The expected value of transmission time is:
[…]
where $f(t)$ is the probability density function, $f(t)=P^{\prime}(t)=\frac{1}{t_{e}-t_{s}}$.
This gives us $E(t)=\frac{t_{s}+t_{e}}{2}$.
For both cases ($\frac{N}{V}<\frac{Q}{Q_{i}V}$ and $\frac{N}{V}>\frac{Q}{Q_{i}V}$), the expected period for flow $i$ is:
$E_{i}=\frac{Q_{i}N+Q}{2Q_{i}V}$ (6)
Taking the average of expected periods across all flows, we have:
$E$ $=\frac{\sum_{i=1}^{N}Q_{i}\cdot E_{i}}{Q}$ (7)
$=\frac{\sum_{i=1}^{N}Q_{i}\cdot\frac{Q_{i}N+Q}{2Q_{i}V}}{Q}$ (8)
$=\frac{N\sum_{i=1}^{N}Q_{i}+Q\sum_{i=1}^{N}1}{2QV}$ (9)
$=\frac{NQ+QN}{2QV}$ (10)
$=\frac{N}{V}$ (11)
This result confirms that our model achieves fair token resource allocation on average, balancing the needs of both high-rate and low-rate flows while preventing FPGA overload, which aligns with our design objectives as outlined in the main text.